OpenStudy (anonymous):

Points V, W, X, Y, and Z are graphed on a number line. The coordinate of point V is –8. What points are at a distance of 5 units from point V? Choose all answers that are correct. A. W = –13 B. X = –8 C. Y = –3 D. Z = 3

OpenStudy (anonymous):

-8 +5 or -8-5 -8+5=-3 -8-5=-13

OpenStudy (anonymous):

so a b and c?

OpenStudy (anonymous):

5 units i would say A

OpenStudy (anonymous):

there could be two answers it says choose all that are correct

OpenStudy (anonymous):

yes it could be A or C..... but the alphabetical order V W X Y Z

OpenStudy (anonymous):

k thanks! :)

OpenStudy (anonymous):

WY